OpenStudy (anonymous):

Given the equation y - 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x - 2y = -12 y = one halfx y = one halfx + 6 y = one halfx + 9

OpenStudy (anonymous):

Given the equation y - 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x - 2y = -12 y = one half x y = one half x + 6 y = one half x + 9